OpenStudy (anonymous):

consider the strophoid r = 2cos(theta)-sec(theta)

OpenStudy (anonymous):

a) sketch the strophoid from\[-\pi/3\] to \[\pi/3\] how do i sketch this without using graphing calculator...?

OpenStudy (anonymous):

b) find the slope of the tangent to the strophoid in terms of \[\theta\] at what points is the slope 0? where is the slope infinite? for the slope equation, i got \[(\sec^2\theta-2\cos2\theta)/2(\sin2\theta+\tan \theta)\] however i dont know how to continue since i do not know how to solve \[\sec^2\theta-2\cos2\theta\]......

OpenStudy (anonymous):

A plot of interest from Mathematica is attached.

OpenStudy (anonymous):

@robtobey thank you!. but can you please tell me how do to graph it using only pen and paper....i could not find any steps or examples in my textbook....

OpenStudy (anonymous):

Sorry, no. That is what computers are for.

OpenStudy (anonymous):

You have to plot \(r(\theta)=2\cos\theta-\sec\theta\) for \(-\dfrac{\pi}{3}\le\theta\le\dfrac{\pi}{3}\). Pick some values of \(\theta\) that are easy to find the cosine and secant of, such as \(\pm\dfrac{\pi}{6}\), starting at one endpoint.